# MOSAIC: Modeling Social AI for Content Dissemination and Regulation in Multi-Agent Simulations

### v1

@@ Abstract
We present a novel, open-source social network simulation framework MOSAIC where generative language agents predict user behaviors such as liking, sharing, and flagging content. This simulation combines LLM agents with a directed social graph to analyze emergent deception behaviors and gain a better understanding of how users determine the veracity of online social content. By constructing user representations from diverse fine-grained personas, our system enables multi-agent simulations that model content dissemination and engagement dynamics at scale. Within this framework, we evaluate three different content moderation strategies with simulated misinformation dissemination, and we find that they not only mitigate the spread of non-factual content but also increase user engagement. In addition, we analyze the trajectories of popular content in our simulations, and explore whether simulation agents’ articulated reasoning for their social interactions truly aligns with their collective engagement patterns. We open-source our simulation software to encourage further research within AI and social sciences: https://github.com/genglinliu/MOSAIC
@@
Genglin Liu1 Salman Rahman1 Elisa Kreiss1 Marzyeh Ghassemi2 Saadia Gabriel1 1 University of California, Los Angeles  2 MIT CSAIL genglinliu@cs.ucla.edu
@@ 1 Introduction
Recent advances in foundation models have led to the emergence of generative agent-based social simulations, where AI-powered users dynamically engage in posting, sharing, flagging, and commenting on content Yang et al. (2024); Gao et al. (2023); Chen et al. (2024a); Wang et al. (2024); Zhou et al. (2023); Park et al. (2022). Unlike traditional survey methods or classical agent-based modeling, simulations driven by large language models enable agents to interact with the environment and each other naturally through rich, human-like dialogue, closely mirroring authentic social behavior. In this work, we introduce MOSAIC, a novel multi-agent AI social network simulation that models content diffusion, user engagement patterns, and misinformation propagation.
Simulation gives us the power to ask counterfactual questions about the complex world. Among different applications of social simulations, content moderation stands out as a pressing challenge due to the real-world harm caused by mis- or disinformation and online influence operations. Previous research has shown that false information not only spreads more rapidly and deeply than truthful content Vosoughi et al. (2018) but also alters public perception in ways that are difficult to reverse Lewandowsky et al. (2012). Addressing this issue requires effective content moderation strategies that can mitigate harm while preserving user engagement and freedom of expression. We embed three moderation strategies into our simulation environment: (1) community-based fact-checking mimicking X and Meta’s Community Notes, (2) independent fact-checking, and a mix of those. We systematically evaluate the impact of these 3 content moderation strategies on misinformation spread, moderation precision/recall, and user engagement dynamics.
@@
  • We build a novel multi-agent simulation from scratch where LLM-powered users dynamically engage with online content, enabling realistic modeling of social behaviors and content diffusion. Interestingly, we find that agents can accurately model individuals, but they’re better at simulating some (more common) demographic groups than others (Section 2).
  • We conduct a comparative study of third-party, community-based, and hybrid fact-checking approaches, quantifying their effectiveness in mitigating misinformation while preserving engagement. We show that misinformation doesn’t spread as fast in an agent simulation as it is commonly observed in human social media, and content moderation strategies can improve not only fact-checking but also engagement (Section 3).
  • We explore how different content and network properties influence diffusion dynamics, offering insights into engagement patterns and how some content/users end up attracting more attention than others. Surprisingly, we find that agents’ individual verbose reasoning may not really reflect their collective action patterns on a group level (Section 4).
By bridging social science observations, game-theoretic modeling Acemoglu et al. (2023) and LLM-driven modeling, our work demonstrates the potential of generative agent simulations as a tool for studying large-scale online behaviors, testing content moderation strategies, and mitigating misinformation risks in the era of generative AI.
2 Social Network Simulation
@@
Category Significant Differences Non-Significant Differences
Age 25-34 (shares) 18-24, 35-44, 45-54, 55-64, 65-74
Gender Male (likes, shares) Female
Asian (comments)
Our AI-driven social network simulates how content spreads, how users engage, and how misinformation propagates within a directed social graph. As illustrated in Figure 1, at its core, the system simulates a dynamic environment where AI agents interact by following others, posting content, reacting (e.g., liking, sharing, commenting), and reporting misinformation. Each agent operates with a persona generated using a question set inspired by AgentBank Park et al. (2024). The main simulation system tracks the progression of time and the evolving state of the network. It is supported by several key components: a relational database that records all user interactions; a content manager that injects new posts into the network; an analytics module that monitors diffusion patterns and user behavior; and a fact-checking system that evaluates the performance of various content moderation strategies.
@@ 2.1 Simulation Flow
The simulation begins with an initialization phase where the system loads experimental configurations, sets up the database (details in Appendix D), generates an initial user population (more details in Appendix A), and establishes follow relationships. Agents are configured to operate under diverse behavioral traits, reflecting real-world variations in social media engagement. In all of our experiments, agents are driven by gpt-4o Hurst et al. (2024) as the foundation model backbone, unless otherwise specified. We do also implement an option to connect agents with open-weight models through SGLang Zheng et al. (2024) or vLLM’s Kwon et al. (2023) inference engines.
At each time step, news content is introduced based on predefined parameters, with agents dynamically responding to their feeds. Agents can optionally generate posts according to their own interests. However during certain controlled experiments we configure them to only engage through reactions such as liking, sharing, commenting, or reporting misinformation. We describe a more general action space and more details of their decision-making process in Appendix E. The visibility of posts evolves based on engagement metrics, simulating algorithmic amplification effects. If fact-checking is enabled, agents incorporate moderation signals, i.e. they are prompted to pay more attention to potentially falsified content or misinformation. adjusting their interactions accordingly (we discuss the content moderation simulation in more depth in Section 3). Throughout this process, the system tracks key statistics, including content reach, user influence, and misinformation spread. At the end of each simulation run, a post-hoc analysis is conducted to assess content diffusion dynamics, user engagement metrics, influence distribution, and the impact of fact-checking interventions. We also keep track of various network properties such as centrality and triadic closures, and perform homophily analysis to examine clustering patterns in user engagement.
@@ 2.2 Human Validation
To validate the veracity of our simulation, we conducted a human study to compare the sharing patterns between humans and LLM agents. We recruited 204 participants via Prolific. [^2] https://www.prolific.com/ More details of our human survey is provided in Appendix B.
@@
In the first phase of this replication study, we conducted a survey to collect demographic data (e.g., age, gender, religion, ethnicity, education level, language, residence, income, political stance) and personal values and behaviors (e.g., time use, priorities, personality of close relationships, social behavior, hobbies, residential history, social goals, meaningful life events, valued friendship traits, financial habits). Inspired by Park et al. (2024), we used this anonymized data to create individualized personas for 204 LLM-driven agents, each corresponding to a human participant.
In the second phase, both participants and their corresponding LLM agents were shown two curated social media snapshots containing 30 posts. They were instructed to respond to each post using a fixed set of actions (e.g., like, dislike, comment, share). The agents, guided solely by their assigned persona profiles, followed the same instructions. We then analyzed and compared engagement patterns between humans and agents, both overall and across demographic groups, to assess how well LLMs can emulate human social media behavior based on persona information alone.
@@
Our analysis compared engagement behaviors between 204 human participants and the same number of persona-replicated AI agents across 30 diverse social media posts, using independent two-sample t-tests for each engagement type. 10% of the articles are false news articles verified by NewsGuard. As illustrated in Figure 2, no statistically significant differences emerged in likes (t = 1.33, p = 0.19) or comments (t = -1.05, p = 0.30), though humans gave slightly more likes (+2.17 per post), and agents posted more comments (+1.87 per post). A marginally significant difference was observed in shares (t = 2.11, p = 0.04), with humans sharing slightly more (+0.80 per post). These results indicate that persona-driven AI agents display engagement patterns that closely mirror those of humans, supporting the realism of our simulation. Further demographic-level analysis (Table 1) found that out of 52 examined demographic subgroups, only 14 showed statistically significant differences (p < 0.05) in at least one engagement metric. Notable discrepancies appeared in the 25–34 age group (shares) and several religious, ethnic, educational, income, and political categories. However, most demographic groups exhibited no significant differences, suggesting that agents simulate typical engagement behavior more accurately for demographics more prevalent in LLM training data. We provide more details about the per-demographic engagement pattern alignment in Appendix B.
Figure 3: Effectiveness of content moderation approaches in promoting factual content. Positive values: factual content receives more engagement. Negative values: misinformation receives more engagement.
@@ 3 Content Moderation in Simulated Social Environment
Method Total Posts News Posts Factual News Misinfo News Precision Recall F1 Score
No Fact-Checking 2878 500 450 50 - - -
Community-Based 1269 490 441 49 0.462 0.490 0.475
Third-Party 1233 450 405 45 0.219 0.156 0.182
Hybrid 809 500 450 50 0.625 0.600 0.612
Table 2: Comparison of different fact-checking methods including dataset size, news distribution, and evaluation metrics.
Figure 4: Comparison of total interactions over time across different fact-checking strategies.
@@ 3.1 Setup
We obtained a data license from NewsGuard [^3] https://www.newsguardtech.com/ to access proprietary information on widespread misinformation narratives tracked by their independent team of journalists. We collected 1,353 examples of false news from their database with release dates up to December 19th, 2024.444We removed non-English articles and de-deduplicated.
To collect real news, we utilize a news aggregation API  [^5] https://newsapi.org/ to retrieve articles published daily from January 31 to February 28, 2025. The system queries the API for all available topics, prioritizing popular articles in English. For each day in the specified range, we extract key information from the retrieved articles, including their title, description, main content, and publication date. Using the NewsAPI, we scraped a total number of 2470 pieces of real news. Since most of them are nonpolitical news from major media outlets, we consider them as "real news" as opposed to the verified political misinformation from the NewsGuard API.
@@
Our simulations involve agentic users interacting with news posts under four different fact-checking conditions: (1) No Fact-Checking, (2) Community-Based Fact-Checking, (3) Third-Party Fact-Checking with an offline LLM that uses its own parametric knowledge, and (4) Hybrid Fact-Checking, where the latter integrates both community-based and third-party verification mechanisms. The simulations starts with 50 agents, and spans over 40 time steps, with agents making interaction decisions based on the perceived veracity of posts and the presence (or absence) of content moderation. At each new step, we randomly introduce up to 2 more agents into the environment to simulate the regular user growth of the social media platform. We analyze both the overall engagement with posts and the effectiveness of fact-checking strategies in suppressing misinformation.
The action space of agents varies across different fact-checking conditions, reflecting different levels of scrutiny and intervention in their social media interactions. In the no fact-checking setting, agents interact freely with the feed, engaging with posts based solely on their interests and beliefs. They can like, share, comment (within a 250-character limit), or ignore posts, without any explicit instructions to assess the accuracy of the content. In the third-party fact-checking condition, the action space remains the same, but the environment implicitly assumes the presence of external fact-checkers who may influence the visibility or credibility of posts. However, the agents themselves do not perform any direct verification. In contrast, the community fact-checking setting expands the action space by allowing agents to add community notes to posts they deem misleading or in need of additional context, as well as rate existing community notes as either helpful or unhelpful. This introduces a participatory element, encouraging agents to contribute to a crowdsourced verification system. Finally, the hybrid fact-checking condition combines elements from both third-party and community-driven verification. Agents can engage with posts as in previous settings while also considering official fact-checks alongside community notes, contributing their own notes and rating those written by others. Across all conditions, agents must select from predefined valid actions, ensuring consistency in response formats. Additionally, when reasoning is enabled, agents are required to justify their interactions by providing a brief explanation for each chosen action, further enhancing the interpretability of their behavior.
@@
The fact checker presented in this code is an automated content verification system designed to identify and address misinformation on a social platform. It works by prioritizing posts for review based on engagement metrics (likes, shares, comments), news classification, and user flags, with special priority given to content that has received community notes in hybrid fact-checking scenarios. The system leverages an LLM (gpt-4o in our experiments) to analyze post content and render verdicts categorized as "true," "false," or "unverified," each accompanied by an explanation, confidence score, and supporting sources from the model’s training data. When posts are deemed false with high confidence ($\geq 0.9$ in standard mode, or $\geq 0.7$ in hybrid mode with community notes), the system automatically takes them down and records the justification. All fact-check results are stored in the database that maintains an audit trail of verdicts alongside ground truth data when available.
@@
We initialize a scale-free network of LLM-powered agents interacting within a directed social graph, using a Barabási–Albert model Barabási and Albert (1999). Misinformation and factual content are injected into the system at controlled rates, with agents dynamically engaging based on their personas and decision-making processes. Each moderation strategy is implemented in a separate experiment, allowing for comparative analysis. We share more details about the experiment configurations in Appendix F.
3.2 False News Do Not Spread Faster than Real News with Simulation Agents
A key insight from our simulation contradicts established results from human social networks: false news does not spread faster than real news (Figure 4). Prior studies on human social behavior have consistently demonstrated that misinformation propagates more rapidly and deeply than factual content (Vosoughi et al., 2018; Zhao et al., 2020). However, in our agent-based simulation, engagement (particularly with sharing) with misinformation does not surpass that of factual news, even in the absence of fact-checking.
From Figure 4, across all experimental conditions, factual news generally maintains higher levels of engagement compared to misinformation. This trend is particularly pronounced in the Third-Party and Hybrid fact-checking setups, where the separation between factual and false news interactions is maximized. Even in the No Fact-Checking scenario, misinformation fails to gain a dominant foothold, suggesting that LLM-driven agents may inherently avoid interacting with unverified or misleading content.

### v2

@@ Abstract
We present a novel, open-source social network simulation framework MOSAIC where generative language agents predict user behaviors such as liking, sharing, and flagging content. This simulation combines LLM agents with a directed social graph to analyze emergent deception behaviors and gain a better understanding of how users determine the veracity of online social content. By constructing user representations from diverse fine-grained actual user personas, our system enables multi-agent simulations that model content dissemination and engagement dynamics at scale. Within this framework, we evaluate three different content moderation strategies with simulated misinformation dissemination, and we find that they not only mitigate the spread of non-factual content but also increase user engagement. In addition, we analyze the trajectories of popular content in our simulations, and explore whether simulation agents’ articulated reasoning for their social interactions truly aligns with their collective engagement patterns. We open-source our simulation software to encourage further research within AI and social sciences: https://github.com/AnonymousPaperRepo/MOSAIC-submission
@@
Genglin Liu1  Vivian Le1  Salman Rahman1 Elisa Kreiss1  Marzyeh Ghassemi2  Saadia Gabriel1 1 University of California, Los Angeles  2 MIT CSAIL genglinliu@cs.ucla.edu
@@ 1 Introduction
Recent advances in foundation models have led to the emergence of social simulations with generative agents, where AI-powered users dynamically engage in social behaviors Yang et al. (2024); Gao et al. (2023); Chen et al. (2024a). Unlike traditional survey methods or classical agent-based modeling, simulations driven by LLMs enable agents to interact with the environment and each other naturally through rich, human-like dialogue, closely mirroring authentic social behavior Wang et al. (2024); Zhou et al. (2023); Park et al. (2022). In this work, we introduce MOSAIC, a novel multi-agent AI social network simulation that models content diffusion, user engagement patterns, and misinformation propagation.
Among different applications of social simulations, content moderation stands out as a pressing challenge due to the real-world harm caused by mis- or disinformation and online influence operations (Jhaver et al., 2023; Young, 2022). Previous research has shown that false information not only spreads more rapidly and deeply than truthful content Vosoughi et al. (2018) but also alters public perception in ways that are difficult to reverse Lewandowsky et al. (2012). Addressing this issue requires effective content moderation strategies that can mitigate harm while preserving user engagement and freedom of expression. We embed three moderation strategies into our simulation environment: (1) community-based fact-checking mimicking X and Meta’s Community Notes, (2) independent fact-checking, and (3) a hybrid approach mixing (1) and (2). We systematically evaluate the impact of these 3 content moderation strategies (along with a baseline of no fact-checking) on misinformation spread, moderation precision/recall, and user engagement dynamics.
@@
  • We build a multi-agent social simulation that is shown to be high fidelity through validation against known social media behavioral phenomena and direct comparison with actual online users. Notably, we find that generative agents are capable of accurately modeling individual engagement patterns, given fine-grained and realistic demographic portraits from user surveys2.2).
  • We conduct a comparative study of third-party, community-based, and hybrid fact-checking approaches, quantifying their effectiveness in mitigating misinformation while preserving engagement. We show that misinformation doesn’t spread as fast in an agent simulation as is commonly observed in human social media, and content moderation strategies can improve not only fact-checking but also engagement3.2, 3.3).
  • We explore how different content and network properties influence diffusion dynamics, offering insights into engagement patterns and how some content/users end up attracting more attention than others. We probe whether the agents’ personas and content topics correlate with engagement. Surprisingly, we find that agents’ individual verbose reasoning may not faithfully reflect their collective action patterns on a group level3.5).
By bridging social science observations, game-theoretic modeling Acemoglu et al. (2023) and LLM-driven modeling, our work demonstrates the potential of generative agent simulations as a tool for studying large-scale online behaviors.
2 MOSAIC Social Network Simulation
@@
Our AI-driven social network simulates how content spreads, how users engage, and how misinformation propagates within a directed social graph. As illustrated in Fig. 1, at its core, the system simulates a dynamic environment where AI agents interact by following others, posting content, reacting (e.g., liking, sharing, commenting), and reporting misinformation. Each agent operates with a persona generated using a question set inspired by AgentBank Park et al. (2024). The main simulation system tracks the progression of time and the evolving state of the network. It is supported by several key components: a relational database that records all user interactions; a content manager that injects new posts into the network; an analytics module that monitors diffusion patterns and user behavior; and a fact-checking system that evaluates the performance of various content moderation strategies.
@@ 2.1 Simulation Flow
The simulation begins with an initialization phase where the system loads experimental configurations, sets up the database (details in Appendix F), generates an initial user population (more details in Appendix B), and establishes follow relationships. Agents are configured to operate under diverse behavioral traits, reflecting real-world variations in social media engagement. In all of our experiments, agents are driven by gpt-4o Hurst et al. (2024) as the foundation model backbone, unless otherwise specified. We do also implement an option to connect agents with open-weight models through SGLang Zheng et al. (2024) or vLLM’s Kwon et al. (2023) inference engines.
At each time step, news content is introduced based on predefined parameters, with agents dynamically responding to their feeds. Agents can optionally generate posts according to their own interests. However during certain controlled experiments we configure them to only engage through reactions such as liking, sharing, commenting, or reporting misinformation. We describe a more general action space and more details of their decision-making process in Appendix G. The visibility of posts evolves based on engagement metrics, simulating algorithmic amplification effects. If fact-checking is enabled, agents incorporate moderation signals, e.g. they are prompted to pay more attention to potentially falsified content or misinformation. We discuss the content moderation simulation in more depth in Section 3. Throughout a simulation run, the system tracks key statistics, including content reach, user influence, and misinformation spread. At the end of each simulation run, a post-hoc analysis is conducted to assess content diffusion dynamics, user engagement metrics, influence distribution, and the impact of fact-checking interventions. We also keep track of various network properties such as centrality [^2] The degree to which an user is central to the network, having outsized or undersized influence. and triadic closures, [^3] A common social phenomenon where users with a mutual connection are more likely to connect to each other. and perform homophily [^4] Another common social phenomenon where similar users are more likely to connect to each other over dissimilar ones. analysis to examine clustering patterns in user engagement (see Appdendix A for an extended technical description of how we define and compute these network metrics). Our focus on these three properties is motivated by a large body of prior work in network theory and the social sciences Rapoport (1953); Granovetter (1973); McPherson et al. (2001); Abebe et al. (2022); Chang et al. (2025).
@@ 2.2 Human Validation
To validate the veracity of our simulation, we conducted a human study to compare the sharing patterns between humans and LLM agents. We recruited 204 participants via Prolific. [^5] https://www.prolific.com/ More details of our human survey are provided in Appendix D.
@@
In the first phase of this replication study, we conducted a survey to collect demographic data (e.g., age, gender, religion, ethnicity, education level, language, residence, income, political stance) and personal values and behaviors (e.g., hobbies, residential history, social goals, meaningful life events, valued friendship traits, financial habits). Inspired by Park et al. (2024), we used this anonymized data to create individualized personas for 204 LLM-driven agents, each corresponding to a human participant.
In the second phase, both participants and their corresponding LLM agents were shown two curated social media snapshots containing 30 posts. 10% of the articles are false news articles verified by an independent and non-partisan team of journalists from NewsGuard. [^6] https://www.newsguardtech.com/ Study participants were instructed to respond to each post using a fixed set of actions (e.g., like, dislike, comment, share). The agents, guided solely by their assigned persona profiles, followed the same instructions. We then analyzed and compared engagement patterns between humans and agents, both overall and across demographic groups, to assess how well LLMs can emulate human social media behavior based on persona information alone.
@@
We analyze the engagement behavior alignment between the human participants and the same number of persona-replicated AI agents, using independent two-sample t-tests for each engagement type. As illustrated in Fig. 2, no statistically significant differences emerged in likes (t = 1.33, p = 0.19) or comments (t = -1.05, p = 0.30), though humans gave slightly more likes (+2.17 per post), and agents posted more comments (+1.87 per post). A marginally significant difference was observed in shares (t = 2.11, p = 0.04), with humans sharing slightly more (+0.80 per post). These results indicate that persona-driven AI agents display engagement patterns that closely mirror those of humans, supporting the realism of our simulation. We also provide more details about the per-demographic engagement pattern alignment in Appendix D.
Figure 3: Effectiveness of content moderation approaches in promoting factual content, across models. Positive values: factual content receives more engagement. Negative values: misinformation receives more engagement.
@@ 3 Content Moderation in Simulated Social Environment
Model Fact-Checking Method Post Statistics Factual Engagement (# of) Misinfo Engagement (# of) Fact-Checking Performance (%)
# of Posts Factual Misinfo Shares Likes Comments Shares Likes Comments Precision Recall F1 Score
deepseek-v3 Unmoderated 420 378 42 14 2478 3968 0 32 117 0.0 0.0 0.0
deepseek-v3 Third Party FC 420 378 42 4 5676 8809 0 114 622 50.0 50.0 50.0
deepseek-v3 Community FC 420 378 42 11 1987 7324 0 15 110 63.3 45.2 52.8
deepseek-v3 Hybrid FC 420 378 42 3 3564 9822 0 42 353 100.0 45.2 62.3
gpt-4o-2024-08-06 Unmoderated 500 450 50 84 69 193 2 1 9 0.0 0.0 0.0
gpt-4o-2024-08-06 Third Party (5p, offline) 450 405 45 752 2387 3800 31 88 255 21.9 63.6 32.6
gpt-4o-2024-08-06 Third Party (5p, online) 420 378 42 430 1883 3020 20 73 335 22.2 100.0 36.4
gpt-4o-2024-08-06 Third Party (10p, online) 420 378 42 616 1917 3066 27 50 131 51.4 90.0 65.5
gpt-4o-2024-08-06 Community FC 490 441 49 80 334 731 3 7 25 45.8 44.9 45.4
gpt-4o-2024-08-06 Hybrid FC 500 450 50 302 2115 4706 7 48 194 73.7 56.0 63.6
claude-3.7-sonnet Unmoderated 420 378 42 55 583 482 2 4 32 0.0 0.0 0.0
claude-3.7-sonnet Third Party FC 368 332 36 571 7411 5526 8 61 280 0.0 0.0 0.0
claude-3.7-sonnet Community FC 420 378 42 15 821 659 0 0 0 100.0 50.0 66.7
claude-3.7-sonnet Hybrid FC 420 378 42 284 8320 7684 0 2 10 100.0 81.0 89.5
Table 1: Comparison of fact-checking (FC) strategies across deepseek-v3, gpt-4o-2024-08-06, and claude-3.7-sonnet under four settings: Unmoderated, Community FC, Third Party FC, and Hybrid FC. For gpt-4o-2024-08-06, third-party setups vary by number of posts reviewed per step (5 or 10) and whether web search is used (online vs. offline). In all other settings, third party is defaulted to checking 5 posts per time step, without web search. Columns are grouped into (1) Post Statistics, (2) Engagement Metrics (shares, likes, comments for factual and misinformation), and (3) Fact-Checking Performance (precision, recall, F1 score on misinformation).
Figure 4: A consolidated view of content engagement and fact-checking metrics, averaged across all models used in our experiments. The first four panels display engagement metrics—specifically, the sum of likes, comments, and shares—under each fact-checking condition. The last two panels show fact-checking metrics, which combine both the number of community notes and note ratings, for the two methods where these are applicable (Hybrid and Community-Based). All values represent the average behavior across models, providing a holistic summary of the system’s dynamics under each experimental setting.
@@ 3.1 Setup
We obtained a data license from NewsGuard to access proprietary information on widespread misinformation narratives tracked by their independent team of journalists. We collected 1,353 examples of false news from their database with release dates up to December 19th, 2024.777We removed non-English articles and de-deduplicated.
To collect factual news articles, we utilize a news aggregation API [^8] https://newsapi.org/ to retrieve articles published daily in legitimate sources from January 31 to February 28, 2025. The system queries the API for all available topics, prioritizing non-political popular articles in English. For each date in the range, we extract key information from the retrieved articles, including their title, description and main content. Using the NewsAPI, we scraped a total number of 2470 pieces of news articles from major media outlets.
@@
Our simulations involve agentic users interacting with news posts under four different fact-checking conditions: (1) No Fact-Checking, (2) Community-Based Fact-Checking, (3) Third-Party Fact-Checking with an independent LLM that uses its own parametric knowledge, and (4) Hybrid Fact-Checking, which integrates both community-based and third-party verification mechanisms. The simulations starts with 50 agents, and spans over 40 time steps, with agents making interaction decisions based on the perceived veracity of posts and the presence of content moderation. At each new step, we randomly introduce up to 2 more agents into the environment to simulate the regular user growth of the social media platform. We analyze both the overall engagement with posts and the effectiveness of fact-checking strategies in suppressing misinformation.
Fact-Checking Settings
The action space of agents varies across different fact-checking conditions, reflecting different levels of scrutiny and intervention in their social media interactions. In the no fact-checking setting, agents interact freely with the feed, engaging with posts based solely on their interests and beliefs. They can like, share, comment, or ignore posts, without any explicit instructions to assess the accuracy of the content. Under this setting, a post is considered “fact-checked" when it has at least one community note with a helpful rating. In the third-party fact-checking condition, the action space remains the same, but the environment implicitly assumes the presence of external fact-checkers who may influence the visibility or credibility of posts. However, the agents themselves do not perform any direct verification. In contrast, the community fact-checking setting expands the action space by allowing agents to add community notes to posts they deem misleading or in need of additional context, as well as rate existing community notes as either helpful or unhelpful. This introduces a participatory element, encouraging agents to contribute to a crowdsourced verification system. Finally, the hybrid fact-checking condition combines elements from both third-party and community-driven verification. Agents can engage with posts as in previous settings while also considering official fact-checks alongside community notes, contributing their own notes and rating those written by others. Across all conditions, agents must select from predefined valid actions, ensuring consistency in response formats. Additionally, when reasoning is enabled, agents are required to justify their interactions by providing a brief explanation for each chosen action, further enhancing the interpretability of their behavior.
@@
The third-party fact checker is represented by an automated content verification system designed to identify and address misinformation on a social platform. It works by prioritizing posts for review based on engagement metrics (likes, shares, comments), news classification, and user flags, with special priority given to content that has received community notes in hybrid fact-checking scenarios. The system leverages an independent LLM instance: gpt-4o (without web browsing capability) and perplexity-sonar-pro [^9] Perplexity AI: https://sonar.perplexity.ai/ (with web browsing) in our experiments. The system categorizes posts as "true," "false," or "unverified," each accompanied by a verdict explanation, confidence score, and supporting evidence sources. When posts are deemed false with high confidence ($\geq 0.9$ in third-party fact-checking mode, or $\geq 0.7$ in hybrid mode with community notes), the system automatically takes them down and records the justification. All fact-check results are stored in the database that maintains an audit trail of verdicts alongside ground truth data when available.
@@
We initialize a scale-free network of LLM-powered agents interacting within a directed social graph, using a Barabási–Albert model Barabási and Albert (1999). Misinformation and factual content are injected into the system at controlled rates. Each moderation strategy is implemented in a separate experiment, allowing for comparative analysis. We share more details about the experiment configurations in Appendix H.
3.2 False News Does Not Spread Faster than Real News with Simulation Agents
A key insight from our simulation contradicts established results from human social networks: false news does not spread faster than real news. Prior studies on human social behavior have consistently demonstrated that misinformation propagates more rapidly and deeply than factual content (Vosoughi et al., 2018; Zhao et al., 2020). However, in our agent-based simulation, engagement (particularly with sharing) with misinformation does not surpass that of factual news.
As shown in Fig. 4, factual news maintains consistently higher engagement levels than misinformation across all four settings. The gap is most pronounced under the Third-Party and Hybrid fact-checking conditions, where factual interactions climb steadily while misinfo remains low. Notably, in the No Fact-Checking scenario, false news still fails to gain a foothold, suggesting that these agents are inherently less likely to propagate unverified or misleading posts. We believe this behavior arises from two factors. First, the agents may rely on their internal confidence to guide sharing decisions, so low-confidence (often false) content is passed over more often. And second, commercial LLMs have been post-trained with safety and helpfulness objectives. This training bias would make our agents less inclined to produce or circulate uncertain or harmful content.
This result highlights a key difference between LLM agents and human users. Human networks amplify sensational or controversial content, but our agents default to a more conservative sharing policy. Future work can explore more sophisticated modeling of cognitive biases or varied trust profiles to better mimic human spread patterns.